OpenStudy (gabylovesyou):

Based on the table of values below, find the slope between points where x = 2 and where x = 6. x y 2 12 4 10 6 4 -3 -2 -1 -1/2

OpenStudy (gabylovesyou):

i think its -1

OpenStudy (gabylovesyou):

@iGreen.

OpenStudy (anonymous):

Slope formula. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{4-12}{6-2}\) Subtract & Divide

OpenStudy (anonymous):

-1 is incorrect.

OpenStudy (alexandervonhumboldt2):

(12-4)/(2-6)

OpenStudy (gabylovesyou):

why subtract 4 - 12 ?

OpenStudy (anonymous):

(2, 12) (6, 4) y2 = 4 y1 = 12 So we subtract 4 - 12

OpenStudy (alexandervonhumboldt2):

slope formula (y1-y2)/(x1-x2)=(12-4)/(2-6)=

OpenStudy (anonymous):

Actually it's y2-y1/x2-x1, but it doesn't really matter..

OpenStudy (gabylovesyou):

OHHH so the answer is -2

OpenStudy (gabylovesyou):

Thanks

OpenStudy (gabylovesyou):

right ? @iGreen.

OpenStudy (anonymous):

Yep, you got it.